Floyd Industries stock has a beta of 1.20. The company just paid a dividend of $.50, and the dividends are expected to grow at 6 percent per year. The expected return on the market is 11 percent, and Treasury bills are yielding 5.9 percent. The most recent stock price for the company is $76.

a.

Calculate the cost of equity using the DDM method. (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)

  DDM method %
b.

Calculate the cost of equity using the SML method. (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)

  SML method %
Answer #1

a.

Calculating Cost of Equity using DDM Model,

Stock Price = D0(1 + g)/(r -g)

76 = 0.50(1.06)/(r - 0.06)

r = 0.50(1.06)/76 + 0.06

r = 0.0670

Cost of Equity = 6.70%

b.

Using CAPM model,

Cost of Equity = Rf + beta(Rm - Rf)

Cost of Equity = 0.059 + 1.20(0.11 - 0.059)

Cost of Equity = 12.02%